OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? ________________________ x 1 2 3 4 y 1 −2 −5 −8 ________________________ y = ? x + ?

OpenStudy (anonymous):

hi again

OpenStudy (anonymous):

@KamiBug

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

so slope is (-2+5)/(1-1)=undefied slope thats no a function x=1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@oncebeaten are you familiar with the slope formula?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

That is slope intercept form. Pick 2 coordinates you want to use to find the slope

OpenStudy (anonymous):

(1,1) and (4,-8)

OpenStudy (anonymous):

wait do you mean n=y2-y1/x2-x1

OpenStudy (anonymous):

Ok, so we will use the formula \(\sf\dfrac{y_2-y_1}{x_2-x_1}\). The point 1, 1 is going to be \(\sf x_1,~y_1\). The point 4, -1 is going to be \(\sf x_2.~y_2\). Following me so far?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ok, so can you substitute those numbers into the formula?

OpenStudy (anonymous):

yea one sec

OpenStudy (anonymous):

1-8/1-4

OpenStudy (anonymous):

Ok, no problem. Take your time

OpenStudy (anonymous):

Almost, and my mistake, I had -1 for \(\sf y_2\). It should have been -8. \(\sf\dfrac{-8-1}{4-1}\)

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

It's fine ^_^

OpenStudy (anonymous):

ok so what now?

OpenStudy (anonymous):

Ok so the slope is -9/3 which is equal to -3 right? Now we need to find the y intercept

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So, to do that, you are going to take 1, 1 and add 3 to the y coordinate and subtract 1 from the x coordinate. Since we are going backwards on the x axis, we need to reverse the slope

OpenStudy (anonymous):

so it would be -4?

OpenStudy (anonymous):

Almost, the y-intercept would be 0, 4

OpenStudy (anonymous):

so just 4 ok

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Now, since slope-intercept form is y=mx+b where m is the slope and b is the y-intercept, can you plug the numbers in?

OpenStudy (anonymous):

y=-3x+4

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

\(\LARGE\cal\color{#ff0074}{No~Problem!}\\\bbox [10pt, #00ffab,border:5pt solid blue ]{\color{#872AF0}{\huge\cal ~\heartsuit user2\heartsuit}}\)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Did that all make sense to you?

OpenStudy (anonymous):

yeah after i read it a couple times. :)

OpenStudy (anonymous):

Ok, just making sure ^_^

OpenStudy (anonymous):

If you have any more questions, feel free to tag me.

OpenStudy (anonymous):

how did you do the fraction signs i know it has to do something with alt and shift

OpenStudy (anonymous):

I will

OpenStudy (anonymous):

`\(\sf\dfrac{-8-1}{4-1}\)` numerator \(\uparrow\) \(\uparrow\) denominator

OpenStudy (anonymous):

does that work on a mac?

OpenStudy (anonymous):

Yep, I am using a mac

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

You can also use the equation button

OpenStudy (anonymous):

what equation button?

OpenStudy (sleepyhead314):

the blue "Equation" button underneath the comment box :)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

http://prntscr.com/5neqh9

OpenStudy (anonymous):

Thanks